The volume of a sphere is given by V=(4/3)πr^3, where r is the radius. Compute the volume of a sphere with the radius of 909pm. State your answer in units of m^3.
1
Expert's answer
2016-08-29T04:55:03-0400
Solution: 1(pm) == 1(picometer) = 1*10-12(m) = 0.000000000001(m)

909pm = 909*10-12(m) = 9.09*10-10(m)
volume of a sphere is given by V=(4/3)πr^3
Here of V=(4/3)*3.14*(9.09*10-10)^3 = 3.14*10-27(m).

Answer: Volume of the sphere = 3.14*10-27(m)